# Fix-CLIP: Dual-Branch Hierarchical Contrastive Learning via Synthetic Captions for Better Understanding of Long Text

### v1

@@ Abstract
CLIP has shown promising performance across many short-text tasks in a zero-shot manner. However, limited by the input length of the text encoder, CLIP struggles on under-stream tasks with long-text inputs ($>77$ tokens). To remedy this issue, we propose Fix-CLIP which includes three novel modules: (1) A dual-branch training pipeline that aligns short and long texts with masked and raw images respectively, which boosts the long-text representation while preserving the short-text ability. (2) Multiple learnable regional prompts with unidirectional masks in Transformer layers for regional information extraction. (3) A hierarchical feature alignment module in the intermediate encoder layers to promote the consistency of multi-scale features. Furthermore, we collect 30M images and utilize existing MLLMs to synthesize long-text captions for training. Extensive experiments show that Fix-CLIP achieves state-of-the-art performance on both long-text and short-text retrieval benchmarks. For downstream applications, we reveal that Fix-CLIP’s text encoder delivers promising performance in a plug-and-play manner for diffusion models with long-text input.
@@ 1 Introduction
CLIP [39] has garnered significant performance across various open-vocabulary tasks. It is widely used as the backbone in Multi-modality Large Language Model (MLLM) [26, 27, 3, 50] and generative models like Stable Diffusion [41].
The success of CLIP is based on the large-scale web-based image-text pairs, which have extremely short effective text length [57]. In fact, images often require dozens of sentences to describe their content adequately. However, CLIP can not understand long text inputs, which severely limits its application to MLLMs and text-to-image generation models. Recently, PixArt-$\alpha$ [4] uses Flan-T5 as the text encoder to increase the length of input tokens from 77 to 120 and injects the obtained text features into DiT [37] to alleviate the deficiency in long-text understanding. Following Long-CLIP [57], our model achieves stronger performance with an input length of 248.
@@ 3.2 Dual-Branch Training Pipeline
It is essential to design distinct encoding strategies for texts of varying lengths to enhance the expressiveness of long texts while maintaining the feature extraction capabilities for short texts. To achieve this, we retain the parameters of the text Transformer blocks from the pre-trained model and modify the position embeddings. In detail, we inherit all the 77 raw position embeddings $PE$ from the pre-trained model and freeze the parameters for short texts. For long texts, we freeze the first 20 position embeddings because the preceding tokens can summarize the major description of the total texts [57]. Then, we expand the remaining position embeddings (from 21 to 77) through the interpolation method to reach four times of the original length, denoted as:
$PE_{l}=Concat(PE[:20],$ $Intpol(PE[20:],4)),$ (1)
@@
Texts of diverse lengths usually correspond to distinct feature spaces, which require customized image features to match. MAE [17] claims that 75% random masked image retains sufficient semantic information. Therefore, aligning random masked images with short captions is an efficient and low-cost pipeline. Specifically, given the raw image patch embeddings $I\in\mathbb{R}^{N\times D}$, we randomly replace $\alpha\times N$ patch embeddings with learnable parameters initialized by 00 to denote the masked images $I_{m}$, where $\alpha$ is the mask ratio and is set as 0.75 at first. Then, we consider the masked image patches $I_{m}$ and short texts $T_{s}$ as the pairs for contrastive learning. Conversely, the long texts often include specific details retained in the raw images. Therefore, we take the raw image patches $I$ and long texts $T_{l}$ as input pairs.
@@ 4.4 Ablation Study
Ablation on different groups of hierarchical feature alignment. As a more reasonable contrastive learning strategy, hierarchical alignment also demonstrates its effectiveness in long-text tasks. We divide all the Transformer blocks into 6 groups both in the image encoder and the text encoder. As illustrated in Tab. 7, applying the hierarchical contrastive learning from the $4$-th group to the $6$-th group achieves 1.6%/3.6% improvement on the DCI benchmark. The last column shows that the GMF loss steadily decreases as the group depth increases. Moreover, the weight of each GMF loss should increase incrementally, as deeper features are more critical for alignment. Finally, we set the weights for GMF loss as $0.2$, $0.4$, and $0.8$ for $4$-th, $5$-th, and $6$-th groups, respectively.
@@
w/ $[1,\quad 6]$ 64.0 63.9 63.1 46.3 4.48
w/ $[3,\quad 6]$ 64.8 65.4 62.8 46.1 3.49
w/ $[4,\quad 6]$ 65.1 66.7 63.4 46.5 2.65
w/ $[5,\quad 6]$ 64.7 65.6 62.7 46.2 1.71
@@ 5 Conclusion
In this work, we propose Fix-CLIP to improve the long-text understanding capability while preserving the short-text ability of CLIP. Considering the distinct feature spaces between short and long texts, we design a dual-branch training pipeline to align short and long texts with masked and raw images respectively. Then, the learnable regional prompts with unidirectional masks are proposed to extract the local features from patch embeddings. We employ a hierarchical alignment module to establish more precise correspondence between intermediate-level textual and visual representations. To explore the performance limits of our model, we leverage MLLMs to synthesize long captions from 30M images and clean the data for training. Fix-CLIP outperforms prior works on numerous open-vocabulary tasks across various training data scales and serves as an effective backbone for diffusion models.
@@ 8.3 Hyperparameters
9 Analysis of Text-to-Image Generation Examples
In this section, we showcase more text-to-image generation examples in long captions to demonstrate the enhancement in understanding long texts. We replace the original text encoder in the stable-diffusion model with that in Long-CLIP [57] or ours. Then, the reconstructed model would be fed with the long captions in the [13] dataset. Due to the divergence between the original text encoder and our text encoder, the model is restrained to generate coarse images. Therefore, an image-to-image refiner model is utilized subsequently to transfer the coarse images to fine images. The final performance is illustrated in Fig. 8. The result of Long-CLIP [57] has confusion in some details, i.e. the background, the direction, and the position relation. Even hallucinations would occur, such as the airplane equipping four jet engines in the $4$-th case. For the comparison, our model correctly describes the detailed information and performs better.
10 Raw Short Caption versus Synthesis Short Caption
@@
Table 13: Above: ablations on the efficacy of region prompts and masks. “Shared Prompts" refers to all the layers utilizing the same shared prompts, “R2P” and “P2R” denote regional prompts attending to all patch embeddings and vice versa. Bottom: the performance comparison of different position embeddings.
11 Visualization of the Similarity Heatmap
We visualize the heatmap of similarity between image features and text features, and compare our results with those of CLIP [39] and Long-CLIP [57], as shown in Fig. 9. To evaluate the performance on short texts, the prompt is set as "a photo of [CLS]". Fix-CLIP demonstrates superior performance over CLIP [39], accurately identifying instances in the image, as illustrated in Fig. 9(a). For long-text understanding, the prompt consists of a major sentence split from the original long-text captions, enabling a direct comparison with Long-CLIP [57]. The corresponding performance is depicted in Fig. 9(b).

### v2

@@ Abstract
CLIP has shown promising performance across many short-text tasks in a zero-shot manner. However, limited by the input length of the text encoder, CLIP struggles on under-stream tasks with long-text inputs (>77>77> 77 tokens). To remedy this issue, we propose Fix-CLIP, which includes three novel modules: (1) A dual-branch training pipeline that aligns short and long texts with masked and raw images, respectively, which boosts the long-text representation while preserving the short-text ability. (2) Multiple learnable regional prompts with unidirectional masks in Transformer layers for regional information extraction. (3) A hierarchical feature alignment module in the intermediate encoder layers to promote the consistency of multi-scale features. Furthermore, we collect 30M images and utilize existing MLLMs to synthesize long-text captions for training. Extensive experiments show that Fix-CLIP achieves state-of-the-art performance on both long-text and short-text retrieval benchmarks. For downstream applications, we reveal that Fix-CLIP’s text encoder delivers promising performance in a plug-and-play manner for diffusion models with long-text input. The code is available at https://github.com/bcwang-sjtu/Fix-CLIP.
[…]
@@ 1 Introduction
CLIP [44] has garnered significant performance across various open-vocabulary tasks. It is widely used as the backbone in Multi-modality Large Language Model (MLLM) [29, 30, 3, 54] and generative models [46, 39, 5, 13].
[…]
@@ 3.2 Dual-Branch Training Pipeline
It is essential to design distinct encoding strategies for texts of varying lengths to enhance the expressiveness of long texts while maintaining the feature extraction capabilities for short texts. To achieve this, we follow Long-CLIP [63] to retain the parameters of the text Transformer blocks from the pre-trained model and modify the position embeddings. We inherit 77 raw position embeddings $PE$ from the pre-trained model and freeze the parameters for short texts. For long texts, we freeze the first 20 position embeddings. Then, we expand the remaining position embeddings (from 21 to 77) through the interpolation method to reach four times of the original length, denoted as:
[…]
@@
Texts of diverse lengths usually correspond to distinct feature spaces, which require customized image features to match. MAE [19] claims that 75% random masked image retains sufficient semantic information. Therefore, aligning random masked images with short captions is an efficient and low-cost pipeline. Specifically, given the raw image patch embeddings $I\in\mathbb{R}^{N\times D}$, we randomly replace $\alpha\times N$ patch embeddings with learnable parameters initialized by 0 to denote the masked images $I_{m}$, where $\alpha$ is the mask ratio and is set as 0.75 at first. Then, we consider the masked image patches $I_{m}$ and short texts $T_{s}$ as the pairs for contrastive learning. Conversely, the long texts often include specific details retained in the raw images. Therefore, we take the raw image patches $I$ and long texts $T_{l}$ as input pairs.
@@ 4.4 Ablation Study
Ablation on different groups of hierarchical feature alignment. As a more reasonable contrastive learning strategy, hierarchical alignment also demonstrates its effectiveness in long-text tasks. We divide all the Transformer blocks into 6 groups both in the image encoder and the text encoder. As illustrated in Tab. 7, applying the hierarchical contrastive learning from the 444-th group to the 666-th group achieves 1.6%/3.6% improvement on the DCI benchmark. The last column shows that the GMF loss steadily decreases as the group depth increases. Moreover, the weight of each GMF loss should increase incrementally, as deeper features are more critical for alignment. Finally, we set the weights for GMF loss as 0.20.20.2, 0.40.40.4, and 0.80.80.8 for 444-th, 555-th, and 666-th groups, respectively.
@@
w/ [1,6][1$\quad$ 6][ 1 , 6 ] 64.0 63.9 63.1 46.3 4.48
w/ [3,6][3$\quad$ 6][ 3 , 6 ] 64.8 65.4 62.8 46.1 3.49
w/ [4,6][4$\quad$ 6][ 4 , 6 ] 65.1 66.7 63.4 46.5 2.65
w/ [5,6][5$\quad$ 6][ 5 , 6 ] 64.7 65.6 62.7 46.2 1.71
@@ 5 Conclusion
In this work, we propose Fix-CLIP to improve the long-text understanding capability while preserving the short-text ability of CLIP. Considering the distinct feature spaces between short and long texts, we design a dual-branch training pipeline to align short and long texts with masked and raw images, respectively. Then, the learnable regional prompts with unidirectional masks are proposed to extract the local features from patch embeddings. We employ a hierarchical alignment module to establish more precise correspondence between intermediate-level textual and visual representations. To explore the performance limits of our model, we leverage MLLMs to synthesize long captions from 30M images and clean the data for training. Fix-CLIP outperforms prior works on numerous open-vocabulary tasks across various training data scales and serves as an effective backbone for diffusion models.
@@ 8.3 Hyperparameters
Table 13: Above: ablations on the efficacy of region prompts and masks. “Shared Prompts” refers to all the layers utilizing the same shared prompts, “R2P” and “P2R” denote regional prompts attending to all patch embeddings and vice versa. Bottom: the performance comparison of different position embeddings.
9 Raw Short Caption versus Synthesis Short Caption
10 Visualization of the Effects of Unidirectional Masking and Region Prompts
In Fig. 8, the regional prompts obtain stronger responses in the corresponding local patches. The red boxes visualize how regional prompts incorporate local features, highlighting the role of Unidirectional Mask. Moreover, the heatmap (b) exhibits higher global responses compared to heatmap (a).
Figure 8: Visualization of the Effects of Unidirectional Masking and Region Prompts.
@@ 10 Visualization of the Effects of Unidirectional Masking and Region Prompts
11 Visualization of the Similarity Heatmap
We visualize the heatmap of similarity between image features and text features, and compare our results with those of CLIP [44] and Long-CLIP [63], as shown in Fig. 9. To evaluate the performance on short texts, the prompt is set as ”a photo of [CLS]”. Fix-CLIP demonstrates superior performance over CLIP [44], accurately identifying instances in the image, as illustrated in Fig. 9(a). For long-text understanding, the prompt consists of a major sentence split from the original long-text captions, enabling a direct comparison with Long-CLIP [63]. The corresponding performance is depicted in Fig. 9(b).
12 Analysis of Text-to-Image Generation Examples
In this section, we showcase more text-to-image generation examples in long captions to demonstrate the enhancement in understanding long texts. We replace the original text encoder in the stable-diffusion model with that in Long-CLIP [63] or ours. Then, the reconstructed model would be fed with the long captions in the [15] dataset. Due to the divergence between the original text encoder and our text encoder, the model is restrained to generate coarse images. Therefore, an image-to-image refiner model is utilized subsequently to transfer the coarse images to fine images. The final performance is illustrated in Fig. 10. The result of Long-CLIP [63] has confusion in some details, i.e. the background, the direction, and the position relation. Even hallucinations would occur, such as the airplane equipping four jet engines in the 444-th case. For the comparison, our model correctly describes the detailed information and performs better.